OpenStudy (anonymous):

If x is a real number, for what values of x is the equation (3x-9)/3=x-3 true? A) all values of x B)some values of x C) no values of x D)impossible to determine

OpenStudy (anonymous):

please help(:

OpenStudy (anonymous):

anyone?

OpenStudy (agent0smith):

Do you know how to simplify the equation, or solve for x?

OpenStudy (anonymous):

cross multiply...I just dont know what the description of x would be..like the answer choices

OpenStudy (anonymous):

i got 3x-9=3x-9 so what would that be called?

OpenStudy (agent0smith):

Well both sides of the equation are equal... what does that mean? Try plugging in a few values for x and see what happens..

OpenStudy (anonymous):

which answer would it be?

OpenStudy (anonymous):

all values of x?

OpenStudy (agent0smith):

If both sides of the equation are equal, what happens when you plug in any value of x? Is the equation always going to be true for any x? If so, then all values of x.

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

Those equations are called identity equations, btw. When you get one side equal to the other like x-3 = x-3, it's an identity equation and must be true for all x.